operating system

Consider a fixed partitioning scheme with equal-size partitions of 216 bytes and a total main memory size of 224 bytes. A process table is maintained that includes apointer to a partition for each resident process. How many bits are required for the pointer?
Answer #1

Size of each partition = 2^16 bytes

Size of main memory = 2^24 bytes

As the partitions are equally sized, the number of partitions in the memory = 2^24 / 2^16

= 2^8 partitions

So the number of bits required for a pointer to identify a partition = 8 bits

answered by: Olili